OpenStudy (anonymous):

Suppose that f(x) = (4x+12)/x. Evaluate the following: f'(-2) = ? f'(5)= ?

OpenStudy (anonymous):

i'm sorry it's ok. i already got the answers.

OpenStudy (anonymous):

i got -3 and -.48 respectively

OpenStudy (nincompoop):

solve for the derivative first use the quotient rule or you can apply sum of divisions \(f(x) = \huge \frac{4x + 12}{x} = \frac{4x}{x} + \frac{12}{x}\)

OpenStudy (nincompoop):

oh okay then close this post

OpenStudy (solomonzelman):

yes, and for f`(2) plug in 2 for x into the derivative and f`(5) plug in 5 for x into the derivative

OpenStudy (solomonzelman):

this would be the slope of the function at x=2, and the slope of the function at x=5.

OpenStudy (solomonzelman):

yes, separation of those fraction is what comes to mind. NOT using the qutient rule of a derivative.

OpenStudy (anonymous):

oh i used the quotient rule

OpenStudy (solomonzelman):

and if you every forget the quotient rule, you can always use logarithmic differnetition to come up to it, which I find very cool.

OpenStudy (nincompoop):

ya or just use the power rule and product rule

OpenStudy (solomonzelman):

yes, that too:)

OpenStudy (solomonzelman):

and product can be proven using log. differnetion too

OpenStudy (nincompoop):

\(\large \frac{1}{x} = x^{-1}\) if you paid your due in algebra, everything in calculus 1 is easy as a pie

OpenStudy (solomonzelman):

yes, or even hardest is as a pi.

OpenStudy (solomonzelman):

but harder.... if you learn wl\ell, anything is easy.

OpenStudy (solomonzelman):

wlell xd